A 1200-kg ore cart is rolling at 10.8 m/s across a flat friction-free surface. A crane suddenly drops of ore vertically into the cart. How fast does the cart move just after being loaded with the ore?

4.20 m/s
6.30 m/s
5.70 m/s
3.80 m/s
Answer #1

Solution:

Given:

Mass of cart (m2) = 1200 kg
Initial velocity of cart (v1) = 10.8 m/s


Mass of ore (m2) = ** NOT GIVEN **

But if you have the value of Mass of ore (m2) with you. then use the following method to solve for final velocity:

Let final velocity be: Vf
Here, According to conservation of momentum: m1 v1 = (m1 + m2) Vf
Therefore: Vf  = { m1 v1 }/ (m1 + m2)

Plug in the given values in above equation to get final answer:
Answer #2
M1*V1/M1+M2=1200*10.8/(1200+989)=5.921
answered by: anonymous